Use the method of separation of variables to find all possible product solutions of x2 Uxx + Uyy + 3 x Ux + 2uy +u = 0.
Answer #1

the-o. 2ay + Q: Use the method of separation of variables for find all possible product Solutions of x ²Uxx + Uyy + 3x Ux + AWhich as Now from x² X + 3 x x + (1-0) X = 0. a Second Order linear differential equation Its AE (Auxiliary Equation) is. x²Now, from Y+ 24+ Cy=o. which is a Second Order linear differential equation Its AE (Auxiliary Eduation) is m2 + 2m + c = 0.